UT 20170321 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (Tuson)
Observer AO: GTayler, DMiller (Tucson)
Telescope Operator: GBechetti

Plan for the night

- Continue AO-com LUCI2 with short images all night

- Targets:

- Z CMa: Bin = 1 on detector, lots os short images/dithers in nB-filters

- APM 08279+5255: long-term stability of system

- Strehl-test by Doug

- MRK 231: closing loop on point source with extended emission, large offsets

- M5/92: strehl as fct of distance from RS

Summary

- Bino dropped again as SX acq. camera was not communicating

- Several RunAO failures and poor image quality at DX, stopped commissioning

Details

All times are UT
Files are luci1/2.20170321.NNNN.fits

01:30 Sunset

Weather: Patchy clouds, should disappear, T = 8 deg, Hum = 30%, low wind

Geno is having troubles to wake up the primary @ SX, JHill brought it up

Upon request by Doug we start with a short script in Bino-mode, they believe that the issues
last night were due to NCPA in Bin=2 used (which they have not commissioned!)
nB: without a working AO@SX tonight, there will be no AFC-testing on sky during this run

nB: It is ridiculous that we went on sky last night with a new firmware from Arcetri microgate
untested on sky. This will go back tomorrow allowing using AO@SX as planned for science

02:20 Execute Bino_ZCMa, preset failed on SX acq. camera is not communicating, Bino-stopped!

Geno de-authorizes LUCI1

02:30 Execute L2_ZCMa, RunAO failed, no frames from WFS, Greg resets system,
resent preset, RunAO failed once more, Gain optimization failed (reason unknown), resent preset
...and RunAO failed even one more time, they finally got it running, loop was closed manually

Image quality is not good, much poorer than on LUCI1.

Some Cirrus rolling in, DIMM seeing still between 0.5 and 1".

#0004-0028

03:42 Execute L2_HD54335, loop opened, skip frame

#0029-0034

Given the fact that the IQ is so poor (the AO-team has no idea where this comes from)
and the persistent problems with the failing RunAO we stop AO-com now

Doug & Greg will talk to the Arcetri-team later tonight

The AO-team does a final-quick test at high-elevation - the problems occurred
at elevation below 50deg.

A-star at 73deg:

#0034 No optical gain or NCPA
#0035 with optical gain
#0036 with optical gain and NCPA on
#0037 loops open

It seems that the problem is the flat used for the shell.

Doug & Greg will talk to the Arcetri-team later tonight

-- %USERSIG{JochenHeidt - 2017-03-20}%
Topic revision: r2 - 21 Mar 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback